The following jobs are waiting to be processed at a small machine​ center:

JOB DUE DATE DURATION (DAYS)
10 240 18
20 232 14
30 240 6
40 245 30
50 255 26

All dates are specified as manufacturing planning calendar days. Assume that all jobs arrived on day 200 ​(yesterday) in the order shown but are not scheduled to begin until day 201 (today).

a. Sequence the jobs according to LPT

According to the LPT rule, the sequence of jobs should be _______________________

b. What is the average completion (flow) time?

c. What is the average job lateness?

d. What is the average number of jobs in the system?
Answer #1

a.

Sequence using LPT rule =>40-50-10-20-30

Average flow time 68.4
Average job lateness 29
Average number of jobs in the system 3.638297872

Calculation

A H B с D E F G 1 ЈОВ DUE DATE Start time DURATION (DAYS) End time Flow time lateness 2 40 245 201 30 230 30 0 3 50 255 230 2

Formula

A B D G с Start time 1. JOB DUE DATE 245 255 DURATION (DAYS) 30 2 40 3 50 201 =E2 26 E End time =C2+D2-1 =C3+D3 =C4+D4 =C5+D5